### v1

# T2Vid: Translating Long Text into Multi-Image is the Catalyst for Video-LLMs

Abstract
The success of Multimodal Large Language Models (MLLMs) in the image domain has garnered wide attention from the research community. Drawing on previous successful experiences, researchers have recently explored extending the success to the video understanding realms. Apart from training from scratch, an efficient way is to utilize the pre-trained image-LLMs, leading to two mainstream approaches, i.e. zero-shot inference and further fine-tuning with video data. In this work, our study of these approaches harvests an effective data augmentation method. We first make a deeper inspection of the zero-shot inference way and identify two limitations, i.e. limited generalization and lack of temporal understanding capabilities. Thus, we further investigate the fine-tuning approach and find a low learning efficiency when simply using all the video data samples, which can be attributed to a lack of instruction diversity. Aiming at this issue, we develop a method called T2Vid to synthesize video-like samples to enrich the instruction diversity in the training corpus. Integrating these data enables a simple and efficient training scheme, which achieves performance comparable to or even superior to using full video datasets by training with just 15% the sample size. Meanwhile, we find that the proposed scheme can boost the performance of long video understanding without training with long video samples. We hope our study will spark more thinking about using MLLMs for video understanding and curation of high-quality data. The code is released at https://github.com/xjtupanda/T2Vid.
1 Introduction
The past few years have seen the rapid progress of Multimodal Large Language Models (MLLMs) [40, 24, 1]. Apart from solving traditional vision tasks (such as VQA), these models also excel in following user instructions and generalizing to new tasks. A mainstream paradigm for developing such models takes a two-stage training strategy. The first stage, pretraining, mainly serves to align vision modality with text and inject various kinds of visual knowledge into the model. In this stage, large-scale datasets of text-image pairs are often used, such as LAION [30] and CC [32], comprising a large proportion of the total compute and injecting abundant vision knowledge into models. Some methods also incorporate OCR and detection-related data to improve foundational capabilities [6, 1]. The second stage, instruction fine-tuning, adapts models to accommodate various tasks and helps generalize to new instructions. Training in this stage typically involves instruction data obtained from self-instruction or adaptation of task-specific datasets (e.g. VQA and chart understanding datasets).
Researchers have recently shifted their focus from single-image models to more advanced ones that support video understanding. Borrowing successful experience from developing image models, some video counterparts are typically trained from scratch, following a similar two-stage training paradigm [7, 18]. Apart from this path, some researchers utilize pre-trained image-LLMs instead. Typical approaches include zero-shot inference [16, 35, 12] and further fine-tuning [22, 18, 15, 7]. Nevertheless, a further inspection of both approaches is still lacking, leaving many underlying factors that influence model performance largely unexplored. To unveil these factors, this work investigates these two approaches. Through a chain of explorations, we finally harvest an efficient scheme for fine-tuning image-LLMs for video understanding, dubbed T2Vid.
We start with the probing of zero-shot inference with Image-LLMs. Through quantitative and qualitative evaluation, we find that: (1) Due to a limited effective context window, it is challenging for image-LLMs to generalize to longer videos by simply using more frames; (2) Instead of understanding the temporal correlations among frames, image-LLMs might adopt an “attend-aggregate” pattern for video understanding. Thus, we switch to investigating the fine-tuning approach since further enhancing temporal understanding capabilities can boost the performance of video understanding. Nevertheless, we find a low learning efficiency problem through a simple data scaling experiment. An inspection of data characteristics suggests that this might be due to a lack of instruction diversity in the training corpus. To address this issue, we propose an efficient scheme to synthesize video-like samples. Specifically, we use existing text instruction data whose sample comprises a (long-context, instruction, answer) triplet. The long-context part is split into multiple segments and then further transformed into images, and the instruction and answer stay intact. Processed in this way, the synthetic samples have the same structure as video instruction data and can be incorporated seamlessly into the training corpus. Moreover, the proposed synthetic method can effectively enrich the instruction diversity in an economical way, thus boosting learning efficiency.
Comprehensive experiments demonstrate that our methods can facilitate efficient fine-tuning of image-LLMs for general video understanding and assist models in the comprehension of long videos. We hope that our findings and proposed scheme can motivate more thinking about MLLMs in video understanding tasks and spark the curation of higher-quality video data.
2 Related Work
2.1 Multimodal Large Language Models
The research on Multi-modal large language models (MLLMs) has garnered wide attention from both industry and academia. Stimulated by the extraordinary capabilities that GPT-4 [29] series have shown, researchers have delved into developing open-sourced models that can compete with or surpass this amazing close-sourced product.
Image-LLMs. To develop image-LLMs, the mainstream approach is to build upon powerful pre-trained LLMs and extend LLMs with the capability to perceive and reason with images [24, 1]. Based on a two-stage training recipe, i.e. image-text alignment training and instruction tuning, the developed model can fulfill a wide range of multimodal user queries and present its answers in user-friendly natural language sentences.
Video-LLMs. Following the success of image-LLMs, subsequent endeavors aim to expand the triumph of image comprehension to more intricate video understanding. Works like Video-ChatGPT [27], VTimeLLM [13], PLLaVA [34] and LLaVA-NeXT-Video [44] attempt further fine-tune image-LLMs to enhance video understanding capability. Other research [22, 18, 15, 7] explores training from pre-trained LLM, following the basic alignment-then-finetuning paradigm similar to image-LLM. These approaches usually involve joint training that mixes image and video data in the training corpus. In this study, we build upon pre-trained image-LLMs and enhance video understanding capabilities through fine-tuning.
2.2 Zero-shot Inference for Video Understanding
Apart from further training, an alternative way is to design training-free approaches to perform zero-shot inference with image-LLMs. For example, IG-VLM [16] arranges multiple video frames into a single image grid and designs corresponding prompts to help models understand the grid format. Given a limited total resolution of image input, there is a trade-off between resolutions and the number of sampled video frames. SF-LLaVA [35] scales up the number of input frames by introducing a dual-stream design, where the slow branch utilizes more spatial tokens extracted with a lower frame rate, while the fast branch is the opposite. Free Video-LLM [12] focuses on efficient inference and introduces independent prompt-guided sampling strategies for spatial and temporal dimensions.
However, these methods are generally evaluated on more traditional benchmarks like MSVD-QA [33], TGIF-QA [14] and ActivityNet-QA [41]). These benchmarks are generally domain-specific and focus on certain basic skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage (especially in longer videos) and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning.
Recently, with the rise of benchmarks specifically designed for MLLMs [9, 18, 26], a more in-depth and comprehensive evaluation has become more accessible. Compared to previous traditional benchmarks, these newly developed benchmarks are generally more challenging, often entailing composite skills and a finer-grained understanding of the video (e.g. the plot in the movie or causal relationships between events), and can be much longer in duration (e.g. up to 60 minutes in the Video-MME benchmark). In this work, we identify the potential limitations of zero-shot inference using these newly developed video benchmarks.
3 Model Formulation
We focus on MLLMs of the mainstream token-level fusion architecture [40], which typically comprises a vision encoder, a connector, and an LLM backbone.
Given a user query and visual input, the vision encoder extracts features from the visual input, and the connector further projects the vision features to align with text embedding space. Then, the processed vision features are sent together with the text embeddings of user queries into the LLM backbone to output natural language response in an auto-regressive way:
$p(w_{o}|\textrm{w}_{V},\textrm{w}_{T})\sim\prod_{t=1}^{L}P(w_{t}|w_{<t},\textrm{w}_{V},\textrm{w}_{T})$ (1)
where $\textrm{w}_{o}=\{w_{o,t}\}_{t=1}^{L}$ is the output word token sequence of length $L$, $\textrm{w}_{V}$ represents the visual tokens processed by a vision encoder and a connector, $\textrm{w}_{T}$ is the text embeddings of the user query.
During our exploration, we mainly utilize three image-LLMs, including Mini-InternVL-Chat-4B-V1.5 [6] (termed as InternVL hereafter), MiniCPM-Llama3-8B-V2.5 [37] (termed as MiniCPM-8B hereafter), and Idefics3-8B-Llama3 [17] (termed as Idefics3-8B hereafter). These instruction-tuned models are trained with massive image data and equipped with strong foundational capabilities. To support higher-resolution vision input, these models adopt the patchifying technique [38, 21, 23] with a dynamic resolution scheme, where an image can be cropped into multiple sub-images according to different aspect ratios. Specifically, InternVL supports up to 13 sub-images where each one is converted into 256 visual tokens; MiniCPM-8B slices images into a maximum of 10 patches, where each is represented by 96 visual tokens; Idefics3-8B supports up to 26 sub-images, each comprising 169 tokens.
4 Evaluation Setup
To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [9], MVBench [18], and TempCompass [26] as our benchmarks. We mainly use Video-MME in our probing since it comprehensively covers different visual domains and temporal ranges. We do not use traditional video-QA benchmarks (e.g. MSVD-QA [33], TGIF-QA [14], ActivityNet-QA [41]) since these benchmarks are generally domain-specific and focus on certain skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel. We illustrate more about the benchmarks used as follows:
Video-MME is a comprehensive benchmark designed for the evaluation of video-LLMs. For temporal coverage, videos of short length (up to 2 minutes), medium length (4–15 minutes), and longer duration (30–60 minutes) are included. The videos and annotations are manually collected and filtered. We only use the raw frames without the subtitles to evaluate video understanding capabilities.
MVBench designs a set of 20 video tasks that cover both perception and cognition, such as scene transition and episodic reasoning. Compared to Video-MME, the videos are sourced from existing benchmarks, and the QAs are automatically generated for the 20 pre-defined tasks.
TempCompass focuses on fine-grained temporal aspects, such as action, speed, and attribute change. The videos and metainformation are manually collected, after which annotations are generated by LLMs with the aid of metainformation. We use the multiple-choice QA (MCQ) format to align with other benchmarks.
To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment. More details about the implementation of this evaluation scheme are available in the Appendix.
5 Motivation
5.1 The Limitations of Zero-shot Reasoning
Methods LLM Params (B) Short Video Medium Video Long Video Overall
  Video-LLaVA [22] 7 45.3 38.0 36.2 39.9
VideoChat2 [18] 7 48.3 37.0 33.2 39.5
Chat-UniVi-v1.5 [15] 7 45.7 40.3 35.8 40.6
VideoLLaMA 2 [7] 8 56.0 45.4 42.1 47.9
InternVL (Zero-shot) [6] 3.8 61.3 51.8 44.3 52.5
VITA [10] 8x7 65.9 52.9 48.6 55.8
Kangaroo [25] 8 66.1 55.3 46.6 56.0
  InternVL (200K full data) 3.8 66.7 54.2 48.1 56.3
InternVL (30K hybrid data) 3.8 67.0 53.7 49.3 56.7
Table 1: Accuracy comparisons of different methods on the Video-MME benchmark, ranking in ascending order regarding overall performance. Our methods use only 15% of the total sample size compared to the full volume video datasets (200K) for fine-tuning and achieve comparable performance. 30K hybrid data comprise 20K data sampled from the full video datasets and 10K synthesized from text data using our T2Vid method. Bold digits indicate the best performance. Lightgray denotes zero-shot inference with image-LLM.
5.1.1 Lack of Frame Count Generalization
A look at zero-shot performance. To gain an initial grasp of zero-shot performance, we compare the results with some representative video-LLMs of similar LLM parameter size, including Video-LLaVA [22], VideoChat2 [18], Chat-UniVi-v1.5 [15], VideoLLaMA 2 [7], VITA [10], and Kangaroo [25], as summarized in Tab. 1.
The table shows that, through zero-shot inference, the image-LLM model already outperforms a variety of video-LLMs of larger LLM parameter sizes. This suggests that a well-trained image-LLM can somehow generalize to video understanding. The zero-shot video understanding capability can be attributed to the widely used training technique “patchifying”, which crops a large image into multiple sub-images and exempts the need for large-scale fine-tuning of vision encoders. Processed in this way, an image can be viewed as a series of sub-images, and the model processes all these sub-images together to understand the overall image. Similarly, a video is naturally a sequence of image frames. Due to this similarity in input structure, it is reasonable that image-LLMs somehow generalize to more frames without further fine-tuning.
Figure 1: Zero-shot performance on Video-MME when inference with a different number of frames. Image-LLMs can generalize to input more frames than seen in the training process but to a rather limited extent, making it hard to truly understand longer videos. The consistent increase of input frames brings limited gains to the performance or can even be detrimental. “max training frames” denotes the max number of patches used for a single image when training; “max training context” denotes the maximum context window set when training; “max LLM context” denotes the context window of the original LLM backbone before further multimodal training.
The influence of different frame counts. Now that image-LLMs can achieve competitive performance, we further probe the generalization potential w.r.t. input frames. Specifically, we experiment with different frame counts and visualize the changes in performance, as shown in Fig. 1. In the figure, we can observe that the performance generally first improves as more frames are available (around the green dotted line), then stays relatively stable or gradually drops (before hitting the red dotted line). The phenomenon indicates that though image-LLMs can generalize to more image numbers than seen in the training process, there is potentially an upper limit of frame number that models can effectively process. Therefore, zero-shot inference for longer video understanding can be challenging. More results and analyses in this regard are available in the Appendix.
5.1.2 Lack of Temporal Understanding Capabilities
Figure 2: A qualitative case using the prompt “Describe this video.”, where the video centers around the introduction of a trip to a town. Compared with the fine-tuned model, the zero-shot case reasons with and describes video frames in a relatively independent way without understanding the temporal correlations. This pattern follows the common behavior of the image model. The keyframes are highlighted with colors (except for red, which is used to indicate hallucinations). The numbers on the upper-left corner of each frame indicate frame order. Zoom in for better visibility.
Since image-LLMs are not trained with any video data, it is intriguing to investigate how these models can reason with videos. This prompts us to further probe if the image-LLM really understands a video and the temporal structure or just attends to multiple video frames and reasons with the information extracted from these frames, just like dealing with a patchified image. Our case study indicates that the latter case (which we call “attend-aggregate” pattern) is more likely. Specifically, we sample some videos and prompt the model in a zero-shot way, such as “Describe this video.”. A typical qualitative case, including the answers of zero-shot inference and fine-tuned model, is shown in Fig. 2.
As shown in the figure, though the model answers in a tone as if it is viewing a video, it does not really understand the temporal relationships between frames but instead aggregates information from different frames in a relatively independent way. More specifically, the text in green (corresponding to the 4th and 8th frames) mixes together the subtitle information. However, the two descriptions actually have inherent temporal structure, where the 4-th frame gives an overall introduction (of the town “Hallstatt”) and the 8-th frame introduces specific scenic spots. For the purple part (the 5th frame), the model mistakenly takes the description of transportation in town as the video’s title. Notably, though part of the upper half of the text has not shown up completely, the model successfully infers the complete text using language prior, demonstrating the model’s strong OCR capabilities. Then, in the second paragraph, the model goes back to describe the earlier 2-th and 3-th frame, respectively, using phrases like “in the background” and “another image” as if it is processing multiple images rather than a holistic video. This is expected since, when dealing with a patchified image, the model often attends to multiple sub-images, extracts information from them, and finally assembles a final answer rather than focusing on the temporal structure within these patches.
In contrast, the fine-tuned video model learns more temporal relationships and thus presents a more coherent description of frame sequence, using correct phrases as connections such as “the video starts with” and “transitions to”. Nevertheless, we notice that hallucinations [19, 39] also increase. Besides inherent noise in the training corpus, the phenomenon is highly correlated with the data bias. For example, many answers of samples in the ShareGemini dataset start with phrases like “The video begins with a black screen”, contributing to a strong prior in the fine-tuned model. We list more qualitative cases and analyses in the Appendix.
(a) Share-Gemini
(b) Video-ChatGPT
Figure 3: Video length statistics of ShareGemini and Video-ChatGPT datasets. Both datasets mostly cover videos shorter than 3.5 minutes. We extract video frames at an FPS of 1. For better visibility, we pick samples with frame numbers lower than 99.9 percentile for visualization.
5.2 The Pitfalls in Fine-tuning
In light of the limitations of zero-shot inference with image-LLMs, we turn to investigate the fine-tuning approach for improving the video understanding capabilities of models. In this section, we first illustrate the training settings and then the experimental findings.
5.2.1 Training Datasets
With 8 NVIDIA A800 GPUs, we train end-to-end (except for InternVL-4B, where we freeze the vision encoder) with a learning rate of 5e-6 and a global batch size of 64. During our investigation, we utilize two representative types of datasets, i.e., video-caption pairs and video instruction data. Specifically, we choose the ShareGemini [31] dataset and the Video-ChatGPT [27] dataset as caption and instruction data, respectively. We extract video frames at an FPS of 1 for each video. In consideration of efficiency, we use up to 64 frames for InternVL-4B and 24 frames for both MiniCPM-8B and Idefics3-8B. When the total number of frames exceeds the thresholds, we uniformly downsample the video frames. We show the statistics of video lengths in Fig. 3 and provide a more detailed introduction of the two datasets below.
ShareGemini-Webvid-core100k. It is a video caption dataset with 100K samples in total. The videos are curated from WebVid [2], a web-scale video-caption dataset covering open domains and general subjects. Regarding duration, the dataset mainly contains short videos with lengths shorter than 30 seconds.
The captions are annotated by calling the strong Gemini-1.5-Pro [11] API. To ensure the diversity of video content, an advanced clustering algorithm [3] is used to filter out highly similar videos. For simplicity, we refer to this dataset as ShareGemini in the following parts of the paper.
Video-ChatGPT. The video instruction dataset contains 100K video-instruction pairs. The videos in this collection are derived from ActivityNet [4]. The dataset’s duration is more diverse and uniform, yet the average video length is no more than 3.5 minutes. There are broadly three types of instructions: video summarization, questions about video content, and creative/generative tasks.
The dataset is annotated in a semi-automatic manner. Some of the data samples are manually annotated by human annotators by refining and enriching the video captions. Other instruction data are generated by GPT-3.5 with the aid of off-the-shelf dense prediction and captioning models.
5.2.2 Low Learning Efficiency of Fine-tuning
Setting S M L Overall
Zero-shot 61.3 51.8 44.3 52.5
30K sampled data 66.2(+4.9) 53.3(+1.5) 47.4(+3.1) 55.7(+3.2)
200K full data 66.7(+0.5) 54.2(+0.9) 48.1(+0.7) 56.3(+0.6)
Table 2: Accuracy comparisons of different settings on the Video-MME benchmark, ranking in ascending order regarding overall performance. S, M, and L indicate short, medium, and long videos, respectively. Zero-shot indicates direct inference with the image-LLM, and full data setting indicates fine-tuning with all the video data samples in the training corpus. The performance gains of the latter two settings are calculated relative to the previous row.
We experiment with fine-tuning with only parts of the video data and with the full volume. Specifically, we randomly sample 15% of the total video samples from each video dataset and compare the performance as in Tab. 2. As shown in the table, compared with the short video set, the gains in medium and long video sets are relatively smaller. This can be attributed to the fact that the video lengths in the training corpus are mostly shorter than 3.5 minutes, rendering a large gap in the temporal range between training and testing (in the evaluation phase, video lengths range from 4–15 minutes and 30–60 minutes for medium and long videos, respectively).
More importantly, the additional gain of scaling up training volume from 30K to 200K data samples is limited (+0.6 overall). In view of the huge difference in data samples used, the learning efficiency of fine-tuning with these video datasets can be quite limited. The phenomenon also suggests that there could be high redundancy in the training corpus.
5.2.3 A Look at Instruction Diversity
Figure 4: t-SNE plot of instruction distribution of video datasets - ShareGemini and Video-ChatGPT. 5,000 instructions are sampled from each dataset. ShareGemini exhibits some clear clusters, while Video-ChatGPT is relatively more diverse.
Prior studies have underscored the importance of instruction diversity for fine-tuning LLMs [46] and image-LLMs [42]. These findings prompt us to conduct an inspection of data in this aspect. Thus, we follow previous approaches [36, 45] to visualize the diversity of instructions in the training corpus. Specifically, we randomly sample 5,000 instructions from ShareGemini and Video-ChatGPT, respectively. Then, the instructions are embedded and visualized using the t-SNE technique, as shown in Fig. 4. In the figure, the distribution of ShareGemini exhibits 9 clear clusters. This is because the dataset samples from a fixed pool of 9 templates as instructions, each of which is a variant of “Describe this video in detail.” In contrast, the distribution of Video-ChatGPT is more diverse, as it includes not only video summarization but also questions related to specific video content. Overall, the instructions for these video samples are not diverse enough.
6 Our Methods
Since currently available video data are limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity using synthetic data.
Figure 5: A conceptual illustration of our scheme for synthesizing video-like samples from text data. For each (long-context, instruction, answer) triplet, we split the long context into segments and transcribe these segments into a series of images, simulating the structure of a video. Zoom in for better visibility.
A rich source of instruction data lies in the text domain, and they can effectively enrich the diversity of instruction. Nevertheless, there is inherently a modality gap between the text and visual domain. To better utilize these data, we bridge the modality gap by synthesizing images with the text. Fig. 5 illustrates our overall idea. Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments and embed the text of each segment into a sequence of images. For example, the context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g. an inquiry to summarize the given paragraphs and the corresponding synopsis. After the transformation, the structure of these instruction data is exactly the same as the video counterparts, and we can incorporate the synthetic data into the video training corpus without extra processing.
Figure 6: t-SNE plot of instruction distribution after applying our proposed methods. By incorporating text data, i.e. LongAlpaca [5] and LongQLoRA [5], the overall distribution becomes more diverse. 5,000 instructions are sampled from each dataset.
Our proposed scheme enjoys three benefits: (1) Mixing in text data can effectively enrich the instruction diversity (Fig. 6), thus improving the learning efficiency for video fine-tuning; (2) Images synthesized from text can emulate the 1D temporal structure of video frames since text segments are generally correlated in the context, thus mitigating the gap between common video samples and synthetic ones; (3) Text data are easier to collect than video samples. Thus, utilizing synthetic data can be economical.
On a related note, LLaMA-VID [20] and LongVA [43] also utilize text data to assist video understanding, and we underscore the differences here: (1) Different objectives: LLaMA-VID and LongVA use text data in the fine-tuning and continued pre-training stage, respectively, in order to expand the context window for long video understanding. In contrast, our methods mainly utilize the diversity of instructions in the text data to facilitate efficient fine-tuning. (2) Different processing methods: LLaMA-VID and LongVA directly use the pure text data in training, while our methods transform text samples into video-like samples and unify the training format.
Data Mix S M L Overall
30K Share-Gemini 65.7 52.8 46.1 54.9
30K Video-ChatGPT 66.3 53.0 47.3 55.6
15K Share-Gemini 15K Video-ChatGPT 66.2 53.3 47.4 55.7
10K Share-Gemini 10K Video-ChatGPT 10K synthetic 67.0 53.7 49.3 56.7
10K Share-Gemini 10K Video-ChatGPT 10K pure text 67.3 52.4 47.7 55.8
Zero-shot 61.3 51.8 44.3 52.5
200K full data 66.7 54.2 48.1 56.3
Table 3: Accuracy comparisons of different data compositions on the Video-MME benchmark. Our proposed scheme achieves an overall performance superior to other data mixes of the same amount (30K) and even more data (200K). The Zero-shot and 200K full data fine-tuning settings are listed for comparison.
Video-MME MVBench TempCompass Avg.
Short Medium Long Overall
Text-aided Models
LLaMA-VID [20] 40.6 40.2 34.9 38.6 41.9 45.6 42.0
LongVA [43] 61.1 48.8 45.4 51.8 50.7 56.9 53.1
MiniCPM-8B [37]
Zero-shot 58.3 46.1 40.1 48.2 42.9 49.1 46.7
200K data 60.3 47.3 44.7 50.8 48.0 54.7 51.2
30K data 61.9 51.7 45.6 53.0 48.4 56.8 52.7
Idefics3-8B [17]
Zero-shot 61.3 47.0 45.1 51.2 49.6 55.9 52.2
200K data 62.0 53.1 44.9 53.3 50.7 62.9 55.6
30K data 67.1 53.8 48.0 56.3 51.6 62.3 56.7
Table 4: The performance of different models/settings on three video benchmarks. Our methods use only 15% of the total sample size compared to the full video sets (200K) for fine-tuning and achieve similar or even superior performance. The highest performance of a model among all settings is highlighted in bold. LLaMA-VID and LongVA also use pure text data but without transformation, which is different from our methods and objectives.
7 Evaluation on Proposed Methods
7.1 Ablation on Different Data Compositions
In order to examine the impact of different data compositions and validate the effectiveness of the proposed method, we conduct an ablation study and construct the following settings with the same amount of total data samples:
  • 30K video samples from ShareGemini.
  • 30K video samples from Video-ChatGPT.
  • 15K video samples from ShareGemini and 15K from Video-ChatGPT, respectively.
  • Our proposed scheme: 10K samples each from ShareGemini and Video-ChatGPT, plus 10K samples synthesized from text data (5K from LongAlpaca and 5K from LongQLora, respectively).
  • Same video samples as above (20K in total), plus 10K samples of corresponding pure text data.
The effectiveness of our proposed methods. As shown in Tab. 3, comparing the first three rows, we can find that when using the same amount of video samples, training only with ShareGemini is not as effective as using more diverse data compositions. Meanwhile, using the same amount of data, our proposed scheme (the 4th row) achieves the best performance. Moreover, when compared with the full 200K data fine-tuning setting, our proposed scheme uses much fewer data samples (only 15%) to achieve comparable performance, and the training cost reduces from 276.8 GPU hours to 33.6 GPU hours, making an 8.2x speedup. The overall results demonstrate the importance of instruction diversity and the effectiveness of our proposed methods.
Notably, replacing the synthetic data with the original pure text counterpart achieves an overall inferior performance. We hypothesize that this is due to the inherent domain gap between vision and text. Thus, to simulate the structure of video sequences, transcribing long text into images might be necessary.
Can synthetic data help models understand longer videos? Interestingly, in the training stage, we only utilize synthetic samples of long multimodal context instead of authentic long video samples. However, on the long video benchmark set, our proposed method still achieves a score that is 1.2 points higher than the full data training (as shown in the 4th row compared to 200K full data in Tab. 3). This result suggests that fine-tuning with a long multimodal context can enhance the comprehension of longer videos. In the following section, we will present additional results to further illustrate this point.
7.2 Generalization Ability of Our Proposed Scheme
We further verify the generalization ability of the proposed scheme and evaluate our methods on more image-LLMs of larger parameter sizes, including MiniCPM-8B and Idefics3-8B, across different benchmarks, as shown in Tab. 4.
General effectiveness of the proposed scheme. The results in the table suggest that on both models, though fine-tuning with fewer samples, our proposed scheme achieves performance on par with or better than fine-tuning with the complete video datasets. We also compare with related methods, LLaMA-VID and LongVA, which also utilize text data but in the original form of pure text, in order to expand the context window. In contrast, our methods excel in enriching the instructions via synthesizing video-like data from text data.
Boost in long video understanding. On the long video sets of the Video-MME benchmark, our methods provide consistent gains compared to the full data fine-tuning setting, though we use video samples (most no longer than 3.5 minutes in the training sets) of lengths much shorter than the test sets (30-60 minutes). The results indicate that our proposed methods can assist models in understanding long videos.
8 Conclusion
This paper revisits the two mainstream approaches of utilizing image-LLMs to perform video understanding. A chain of investigations reveals possible drawbacks in both approaches, underscoring the bottlenecks of zero-shot inference and the easily ignored aspects of fine-tuning. Among these discoveries, the limited instruction diversity of current video data is notable, leading to downgraded learning efficiency. Aimed at this specific issue, we develop an efficient method, dubbed T2Vid, to enrich the instructions in the training corpus. Comprehensive evaluations demonstrate the effectiveness of the proposed method. We hope the findings from this paper can spark deeper thinking about using MLLMs for video understanding and how to curate video data of higher quality.
Appendix A Implementation Details
A.1 Answer Judging
We notice that MiniCPM-8B [37] often fails to follow instructions properly when we explicitly ask the model to “Answer with the option’s letter from the given choices directly”, making simple exact matching inaccurate. Specifically, the model often prepends or appends additional text other than the option letters, e.g.Answer: B. Pink.”, or gives additional explanations apart from the answer.
To cope with these issues, we adopt a combination of exact matching and LLM matching for assessment. Specifically, we strip the prefixes such as “Answer:” from the prediction and try to use regular expression matching to find the option letter. When the exact matching scheme fails, we use an LLM (Llama-3.1-8B-Instruct [28]) to find an option closest to the model prediction. When the LLM matching fails, a placeholder outside of the available options (such as “Z”) is returned to denote a wrong answer. Our judging prompt for the LLM is modified from VLMEvalKit [8], as shown in Tab. 5.
Appendix B Additional Results in Experiments
B.1 Results on Different Frame Counts
Figure 7: Accuracy of MiniCPM-8B and Idefics3-8B on the Video-MME benchmark when inference with different numbers of video frames. As input frames increase, the performance of both models stays relatively stable and does not improve much. “max training frames” denotes the max number of patches used for a single image when training; “max training context” denotes the maximum context window set when training; “max LLM context” denotes the context window of the original LLM backbone before further multimodal training.
We present the performance of different input frame numbers of MiniCPM-8B [37] and Idefics3-8B [17] in Fig. 7. As input frames increase, the performance within the context window of the LLM backbone is more stable than the InterVL-4B, which is of a smaller parameter size. Nevertheless, the performance still does not improve much as more frames are available, again indicating that the effective context is limited. Thus, simply expanding the LLM context window for image-LLMs to facilitate long video understanding would not be very practical. As in the main paper, we attribute this phenomenon to the lack of capabilities of image-LLMs to truly understand the temporal relationships among frames. Additional case studies related to this topic are available in the following section.
B.2 Case Study
In the main paper, we identify the drawback of image-LLMs generally failing to understand the temporal relationships among video frames. In this section, we supplement more cases and analyses on this topic.
Figure 8: A case study using the prompt “Describe this video.”, where the video is a documentary of the Arctic environment. Compared with the fine-tuned model, the zero-shot case reasons with and describes video frames without understanding the temporal correlations. This pattern follows the common behavior of the image model. The keyframes are highlighted with colors (except for red, which is used to indicate hallucinations). The numbers on the upper-left corner of each frame indicate frame order. Zoom in for better visibility.
In case 1 (Fig. 8), we list a case where we instruct models to describe the video, while the image-LLM fails to understand the separate scenes and temporal relationships but instead focuses on several specific frames and describes rather independently. In contrast, the model fine-tuned with video data can more accurately distinguish the actions happening in the video and understand the connections in a temporally coherent way.
Figure 9: A case study designed to test the temporal reasoning capabilities of models. In the video, three polar bears show up at the beginning, and in the end, the scene changes into a video shot of two polar bears. The image-LLM fails to recognize the change through zero-shot prompting. Note that we have verified that the image-LLM can correctly recognize the number by inputting a single frame separately.
In case 2 (Fig. 9), we ask the two models to identify the temporal changes in object counts. The fine-tuned model correctly recognizes the changes in the number of polar bears as scenes change in the video clip, while the image-LLM fails to do so and seems to give the answer according to the last few frames. This is expected since image-LLMs are trained with static images rather than dynamic videos, where the latter inject temporal concepts into models.
System message
You are an AI assistant who will help me to match an answer with several options of a single-choice question.
Prompt
You are provided with a question, several options, and an answer, and you need to find which option is most similar to the answer.
If the meaning of all options are significantly different from the answer, output Z. Your should directly output a single uppercase character, such as A, B, C, D (if they are valid options) and Z, and nothing else. Here are two examples.
Example 1:
Question: What is the main object in image?
Options: A. teddy bear.
B. rabbit.
C. cat.
D. dog.
Answer: a cute teddy bear
Output: A
Example 2:
Question: What is the main object in image?
Options: A. teddy bear.
B. rabbit.
C. cat.
D. dog.
Answer: Spider
Output: Z
Now here is the question, options, and the answer, you should match and give me the option letter:
Question: {Question}
Options: {Options}
Answer: {Model Answer}
Output:
Table 5: Template for prompting LLM to perform option matching. {Question} is the specific question of a benchmark sample, and {Options} are corresponding choices of the question. {Model Answer} is the raw prediction of MLLMs.

### v3

@@
-T2Vid: Translating Long Text into Multi-Image is the Catalyst for Video-LLMs
+Sparrow: Data-Efficient Video-LLM with Text-to-Image Augmentation
 Abstract
-The success of Multimodal Large Language Models (MLLMs) in the image domain has garnered wide attention from the research community.
-Drawing on previous successful experiences, researchers have recently explored extending the success to the video understanding realms.
-Apart from training from scratch, an efficient way is to utilize the pre-trained image-LLMs, leading to two mainstream approaches, i.e.
-zero-shot inference and further fine-tuning with video data.
-In this work, our study of these approaches harvests an effective data augmentation method.
-We first make a deeper inspection of the zero-shot inference way and identify two limitations, i.e.
-limited generalization and lack of temporal understanding capabilities.
-Thus, we further investigate the fine-tuning approach and find a low learning efficiency when simply using all the video data samples, which can be attributed to a lack of instruction diversity.
-Aiming at this issue, we develop a method called T2Vid to synthesize video-like samples to enrich the instruction diversity in the training corpus.
-Integrating these data enables a simple and efficient training scheme, which achieves performance comparable to or even superior to using full video datasets by training with just 15% the sample size.
-Meanwhile, we find that the proposed scheme can boost the performance of long video understanding without training with long video samples.
-We hope our study will spark more thinking about using MLLMs for video understanding and curation of high-quality data.
-The code is released at https://github.com/xjtupanda/T2Vid.
+Recent years have witnessed the success of Multimodal Large Language Models (MLLMs) in the vision understanding domain.
+The success of these models can largely be attributed to the dominant scaling law, which states that larger parameter sizes and data volumes contribute to better performance.
+Notably, data scaling has mainly been powered by automatic data pipelines, which center around the self-instruction of LLMs.
+The paradigm has been taken for granted for quite some time, but the study of the effectiveness of scaling with these data has been neglected for a long time.
+In this context, this work revisits scaling with synthetic data and focuses on developing video-LLMs from a data-centric perspective.
+Our main study approach is fine-tuning pre-trained image-LLMs with video data and investigating learning efficiency through data scaling.
+Results from our preliminary experiments reveal a low learning efficiency phenomenon when simply scaling up video data samples, which, through our probing, can be ascribed to a lack of instruction diversity.
+Aiming at this issue, we propose a data augmentation method called Sparrow, which synthesizes video-like samples from pure text instruction data.
+Mixing these synthetic samples with the video data enables a more efficient training scheme.
+Through comprehensive experiments, we demonstrate that our proposed method achieves performance comparable to or even superior to baselines trained with many more samples.
+Meanwhile, we find that incorporating these synthetic samples can boost the performance of long video understanding without training with long video data.
+The code and data examples are available at this link.
+Figure 1: Performance comparison with different schemes of data scaling.
+We introduce Sparrow, a data-efficient training method for video-LLMs that achieves “Train less, gain more” in both general video and long video understanding performance.
+0 sample indicates zero-shot inference with the image-LLM [1].
 1 Introduction
@@
 VQA and chart understanding datasets).
-Researchers have recently shifted their focus from single-image models to more advanced ones that support video understanding.
+Recently, researchers have shifted their focus from single-image models to more advanced ones that support video understanding.
 Borrowing successful experience from developing image models, some video counterparts are typically trained from scratch, following a similar two-stage training paradigm [8, 9].
@@
 Typical approaches include zero-shot inference [10, 11, 12] and further fine-tuning [13, 9, 14, 8].
-Nevertheless, a further inspection of both approaches is still lacking, leaving many underlying factors that influence model performance largely unexplored.
-To unveil these factors, this work investigates these two approaches.
-Through a chain of explorations, we finally harvest an efficient scheme for fine-tuning image-LLMs for video understanding, dubbed T2Vid.
-We start with the probing of zero-shot inference with Image-LLMs.
-Through quantitative and qualitative evaluation, we find that: (1) Due to a limited effective context window, it is challenging for image-LLMs to generalize to longer videos by simply using more frames; (2) Instead of understanding the temporal correlations among frames, image-LLMs might adopt an “attend-aggregate” pattern for video understanding.
-Thus, we switch to investigating the fine-tuning approach since further enhancing temporal understanding capabilities can boost the performance of video understanding.
-Nevertheless, we find a low learning efficiency problem through a simple data scaling experiment.
-An inspection of data characteristics suggests that this might be due to a lack of instruction diversity in the training corpus.
-To address this issue, we propose an efficient scheme to synthesize video-like samples.
+Notably, the success of these models can be largely ascribed to the formidable scaling law, which puts emphasis on scaling up parameter size or data volume for better model performance.
+For the data aspect, the scaling has mainly been driven by automatic data engines, which synthesize massive amounts of data without human labor.
+Nevertheless, the characteristics of learning from these synthesized video data stand out as a critical yet underexplored topic.
+Thus, in this work, we investigate the learning characteristics of video-LLM more deeply from a data scaling perspective.
+Our preliminary data scaling experiments reveal a low data efficiency problem, that is, the performance gains from utilizing multiple times more data is marginal.
+An inspection of data characteristics suggests this might be due to a lack of instruction diversity in the training corpus.
+To address this issue, we propose a data augmentation method, dubbed Sparrow [^1] Inspiration taken from the swiftness of sparrows., to enrich the diversity of instruction.
+The basic idea is to synthesize video-like samples from textual data and mix these synthetic data with the video samples.
 Specifically, we use existing text instruction data whose sample comprises a (long-context, instruction, answer) triplet.
-The long-context part is split into multiple segments and then further transformed into images, and the instruction and answer stay intact.
+The long-context part is split into multiple segments and then further transformed into images, while the instruction and answer stay intact.
 Processed in this way, the synthetic samples have the same structure as video instruction data and can be incorporated seamlessly into the training corpus.
-Moreover, the proposed synthetic method can effectively enrich the instruction diversity in an economical way, thus boosting learning efficiency.
-Comprehensive experiments demonstrate that our methods can facilitate efficient fine-tuning of image-LLMs for general video understanding and assist models in the comprehension of long videos.
-We hope that our findings and proposed scheme can motivate more thinking about MLLMs in video understanding tasks and spark the curation of higher-quality video data.
+Comprehensive experiments demonstrate that our methods can facilitate data-efficient fine-tuning of image-LLMs for general video understanding and assist models in the comprehension of long videos.
+Specifically, using the same number of training samples, our method shows clear advantages over other data schemes.
+It even surpasses the baselines trained with many more samples, achieving high data efficiency (Fig.
+1).
+The contributions of this work include:
+  • We investigate the fine-tuning approach for developing video-LLMs from a data perspective, and shed light on possible factors that lead to low learning efficiency.
+  • We propose a data augmentation method that improves the instruction diversity of training data and facilitates a more efficient training scheme.
+  • We perform comprehensive experiments to evaluate the proposed method and examine its key properties, paving the way for future research in this line.
 2 Related Work
 2.1 Multimodal Large Language Models
-The research on Multi-modal large language models (MLLMs) has garnered wide attention from both industry and academia.
-Stimulated by the extraordinary capabilities that GPT-4 [29] series have shown, researchers have delved into developing open-sourced models that can compete with or surpass this amazing close-sourced product.
 Image-LLMs.
@@
 Video-LLMs.
-Following the success of image-LLMs, subsequent endeavors aim to expand the triumph of image comprehension to more intricate video understanding.
+Following the success of image-LLMs, subsequent endeavors aim to expand the triumph to more intricate video understanding.
 Works like Video-ChatGPT [15], VTimeLLM [16], PLLaVA [17] and LLaVA-NeXT-Video [18] attempt further fine-tune image-LLMs to enhance video understanding capability.
@@
 In this study, we build upon pre-trained image-LLMs and enhance video understanding capabilities through fine-tuning.
-2.2 Zero-shot Inference for Video Understanding
-Apart from further training, an alternative way is to design training-free approaches to perform zero-shot inference with image-LLMs.
-For example, IG-VLM [16] arranges multiple video frames into a single image grid and designs corresponding prompts to help models understand the grid format.
-Given a limited total resolution of image input, there is a trade-off between resolutions and the number of sampled video frames.
-SF-LLaVA [35] scales up the number of input frames by introducing a dual-stream design, where the slow branch utilizes more spatial tokens extracted with a lower frame rate, while the fast branch is the opposite.
-Free Video-LLM [12] focuses on efficient inference and introduces independent prompt-guided sampling strategies for spatial and temporal dimensions.
-However, these methods are generally evaluated on more traditional benchmarks like MSVD-QA [33], TGIF-QA [14] and ActivityNet-QA [41]).
+2.2 Evaluation of Video Understanding
+Early methods [10, 15] are generally evaluated on more traditional benchmarks like MSVD-QA [19], TGIF-QA [20] and ActivityNet-QA [21]).
 These benchmarks are generally domain-specific and focus on certain basic skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage (especially in longer videos) and skill coverage.
 Moreover, the questions asked often involve shallow perception without deeper reasoning.
-Recently, with the rise of benchmarks specifically designed for MLLMs [9, 18, 26], a more in-depth and comprehensive evaluation has become more accessible.
+Recently, with the rise of benchmarks specifically designed for MLLMs [22, 9, 23, 24], a more in-depth and comprehensive evaluation has become more accessible.
 Compared to previous traditional benchmarks, these newly developed benchmarks are generally more challenging, often entailing composite skills and a finer-grained understanding of the video (e.g.
@@
 up to 60 minutes in the Video-MME benchmark).
-In this work, we identify the potential limitations of zero-shot inference using these newly developed video benchmarks.
-3 Model Formulation
-We focus on MLLMs of the mainstream token-level fusion architecture [40], which typically comprises a vision encoder, a connector, and an LLM backbone.
-Given a user query and visual input, the vision encoder extracts features from the visual input, and the connector further projects the vision features to align with text embedding space.
-Then, the processed vision features are sent together with the text embeddings of user queries into the LLM backbone to output natural language response in an auto-regressive way:
-$p(w_{o}|\textrm{w}_{V},\textrm{w}_{T})\sim\prod_{t=1}^{L}P(w_{t}|w_{<t},\textrm{w}_{V},\textrm{w}_{T})$ (1)
-where $\textrm{w}_{o}=\{w_{o,t}\}_{t=1}^{L}$ is the output word token sequence of length $L$, $\textrm{w}_{V}$ represents the visual tokens processed by a vision encoder and a connector, $\textrm{w}_{T}$ is the text embeddings of the user query.
-During our exploration, we mainly utilize three image-LLMs, including Mini-InternVL-Chat-4B-V1.5 [6] (termed as InternVL hereafter), MiniCPM-Llama3-8B-V2.5 [37] (termed as MiniCPM-8B hereafter), and Idefics3-8B-Llama3 [17] (termed as Idefics3-8B hereafter).
+In this work, our study adopts these newly developed video benchmarks.
+2.3 Textual Data for Video Understanding
+Since MLLMs are typically built upon LLMs and thus highly compatible with textual data, some works have explored utilizing pure text data to boost the performance of video understanding.
+Below, we outline the key ideas of these methods and their differences from our method.
+  • Textual data for context expanding.
+Previous works have explored utilizing textual data to expand the context window of base LLMs.
+Specifically, in order to facilitate long video understanding, LLaMA-VID [25] and LongVA [26] incorporate long text data in fine-tuning and continued pre-training stages, respectively, to expand the context window of LLM backbones.
+    Differences: In this work, we (1) adopt textual data as a data augmentation method and (2) use them in the vision form to accommodate the training format.
+  • Synthetic textual data for video understanding.
+This line of work investigates synthesizing textual data that simulates video QA data, aiming to transfer temporal reasoning capabilities from textual training.
+More specifically, TOPA [27] extracts textual captions and object-level information from video frames, while T3 [28] gathers similar information from multiple different images.
+    Differences: These two works seek to boost video understanding with synthetic data, while ours aims to enrich the instruction diversity of the training corpus.
+Moreover, our method does not require calling advanced LLM APIs to build data; instead, our method utilizes existing datasets.
+3 A Probing Study of Data Scaling
+To understand the scaling characteristics of training data, our study starts with fine-tuning with different sample sizes and examines the relationship between training sample size and model performance.
+In this section, we introduce the study’s training and evaluation setup and then illustrate the empirical findings.
+3.1 Training Setup
+3.1.1 Model Setup
+During our exploration, we mainly utilize two image-LLMs, including Mini-InternVL-Chat-4B-V1.5 [7] (termed as InternVL hereafter), MiniCPM-Llama3-8B-V2.5 [1] (termed as MiniCPM-8B hereafter).
 These instruction-tuned models are trained with massive image data and equipped with strong foundational capabilities.
 To support higher-resolution vision input, these models adopt the patchifying technique [29, 30, 31] with a dynamic resolution scheme, where an image can be cropped into multiple sub-images according to different aspect ratios.
-Specifically, InternVL supports up to 13 sub-images where each one is converted into 256 visual tokens; MiniCPM-8B slices images into a maximum of 10 patches, where each is represented by 96 visual tokens; Idefics3-8B supports up to 26 sub-images, each comprising 169 tokens.
-4 Evaluation Setup
-To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [9], MVBench [18], and TempCompass [26] as our benchmarks.
-We mainly use Video-MME in our probing since it comprehensively covers different visual domains and temporal ranges.
-We do not use traditional video-QA benchmarks (e.g.
-MSVD-QA [33], TGIF-QA [14], ActivityNet-QA [41]) since these benchmarks are generally domain-specific and focus on certain skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage and skill coverage.
-Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel.
-We illustrate more about the benchmarks used as follows:
-Video-MME is a comprehensive benchmark designed for the evaluation of video-LLMs.
-For temporal coverage, videos of short length (up to 2 minutes), medium length (4–15 minutes), and longer duration (30–60 minutes) are included.
-The videos and annotations are manually collected and filtered.
-We only use the raw frames without the subtitles to evaluate video understanding capabilities.
-MVBench designs a set of 20 video tasks that cover both perception and cognition, such as scene transition and episodic reasoning.
-Compared to Video-MME, the videos are sourced from existing benchmarks, and the QAs are automatically generated for the 20 pre-defined tasks.
-TempCompass focuses on fine-grained temporal aspects, such as action, speed, and attribute change.
-The videos and metainformation are manually collected, after which annotations are generated by LLMs with the aid of metainformation.
-We use the multiple-choice QA (MCQ) format to align with other benchmarks.
-To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment.
-More details about the implementation of this evaluation scheme are available in the Appendix.
-5 Motivation
-5.1 The Limitations of Zero-shot Reasoning
-Methods LLM Params (B) Short Video Medium Video Long Video Overall
-  Video-LLaVA [22] 7 45.3 38.0 36.2 39.9
-VideoChat2 [18] 7 48.3 37.0 33.2 39.5
-Chat-UniVi-v1.5 [15] 7 45.7 40.3 35.8 40.6
-VideoLLaMA 2 [7] 8 56.0 45.4 42.1 47.9
-InternVL (Zero-shot) [6] 3.8 61.3 51.8 44.3 52.5
-VITA [10] 8x7 65.9 52.9 48.6 55.8
-Kangaroo [25] 8 66.1 55.3 46.6 56.0
-  InternVL (200K full data) 3.8 66.7 54.2 48.1 56.3
-InternVL (30K hybrid data) 3.8 67.0 53.7 49.3 56.7
-Table 1: Accuracy comparisons of different methods on the Video-MME benchmark, ranking in ascending order regarding overall performance.
-Our methods use only 15% of the total sample size compared to the full volume video datasets (200K) for fine-tuning and achieve comparable performance.
-30K hybrid data comprise 20K data sampled from the full video datasets and 10K synthesized from text data using our T2Vid method.
-Bold digits indicate the best performance.
-Lightgray denotes zero-shot inference with image-LLM.
-5.1.1 Lack of Frame Count Generalization
-A look at zero-shot performance.
-To gain an initial grasp of zero-shot performance, we compare the results with some representative video-LLMs of similar LLM parameter size, including Video-LLaVA [22], VideoChat2 [18], Chat-UniVi-v1.5 [15], VideoLLaMA 2 [7], VITA [10], and Kangaroo [25], as summarized in Tab.
-1.
-The table shows that, through zero-shot inference, the image-LLM model already outperforms a variety of video-LLMs of larger LLM parameter sizes.
-This suggests that a well-trained image-LLM can somehow generalize to video understanding.
-The zero-shot video understanding capability can be attributed to the widely used training technique “patchifying”, which crops a large image into multiple sub-images and exempts the need for large-scale fine-tuning of vision encoders.
-Processed in this way, an image can be viewed as a series of sub-images, and the model processes all these sub-images together to understand the overall image.
-Similarly, a video is naturally a sequence of image frames.
-Due to this similarity in input structure, it is reasonable that image-LLMs somehow generalize to more frames without further fine-tuning.
-Figure 1: Zero-shot performance on Video-MME when inference with a different number of frames.
-Image-LLMs can generalize to input more frames than seen in the training process but to a rather limited extent, making it hard to truly understand longer videos.
-The consistent increase of input frames brings limited gains to the performance or can even be detrimental.
-“max training frames” denotes the max number of patches used for a single image when training; “max training context” denotes the maximum context window set when training; “max LLM context” denotes the context window of the original LLM backbone before further multimodal training.
-The influence of different frame counts.
-Now that image-LLMs can achieve competitive performance, we further probe the generalization potential w.r.t.
-input frames.
-Specifically, we experiment with different frame counts and visualize the changes in performance, as shown in Fig.
-1.
-In the figure, we can observe that the performance generally first improves as more frames are available (around the green dotted line), then stays relatively stable or gradually drops (before hitting the red dotted line).
-The phenomenon indicates that though image-LLMs can generalize to more image numbers than seen in the training process, there is potentially an upper limit of frame number that models can effectively process.
-Therefore, zero-shot inference for longer video understanding can be challenging.
-More results and analyses in this regard are available in the Appendix.
-5.1.2 Lack of Temporal Understanding Capabilities
-Figure 2: A qualitative case using the prompt “Describe this video.”, where the video centers around the introduction of a trip to a town.
-Compared with the fine-tuned model, the zero-shot case reasons with and describes video frames in a relatively independent way without understanding the temporal correlations.
-This pattern follows the common behavior of the image model.
-The keyframes are highlighted with colors (except for red, which is used to indicate hallucinations).
-The numbers on the upper-left corner of each frame indicate frame order.
-Zoom in for better visibility.
-Since image-LLMs are not trained with any video data, it is intriguing to investigate how these models can reason with videos.
-This prompts us to further probe if the image-LLM really understands a video and the temporal structure or just attends to multiple video frames and reasons with the information extracted from these frames, just like dealing with a patchified image.
-Our case study indicates that the latter case (which we call “attend-aggregate” pattern) is more likely.
-Specifically, we sample some videos and prompt the model in a zero-shot way, such as “Describe this video.”.
-A typical qualitative case, including the answers of zero-shot inference and fine-tuned model, is shown in Fig.
-2.
-As shown in the figure, though the model answers in a tone as if it is viewing a video, it does not really understand the temporal relationships between frames but instead aggregates information from different frames in a relatively independent way.
-More specifically, the text in green (corresponding to the 4th and 8th frames) mixes together the subtitle information.
-However, the two descriptions actually have inherent temporal structure, where the 4-th frame gives an overall introduction (of the town “Hallstatt”) and the 8-th frame introduces specific scenic spots.
-For the purple part (the 5th frame), the model mistakenly takes the description of transportation in town as the video’s title.
-Notably, though part of the upper half of the text has not shown up completely, the model successfully infers the complete text using language prior, demonstrating the model’s strong OCR capabilities.
-Then, in the second paragraph, the model goes back to describe the earlier 2-th and 3-th frame, respectively, using phrases like “in the background” and “another image” as if it is processing multiple images rather than a holistic video.
-This is expected since, when dealing with a patchified image, the model often attends to multiple sub-images, extracts information from them, and finally assembles a final answer rather than focusing on the temporal structure within these patches.
-In contrast, the fine-tuned video model learns more temporal relationships and thus presents a more coherent description of frame sequence, using correct phrases as connections such as “the video starts with” and “transitions to”.
-Nevertheless, we notice that hallucinations [19, 39] also increase.
-Besides inherent noise in the training corpus, the phenomenon is highly correlated with the data bias.
-For example, many answers of samples in the ShareGemini dataset start with phrases like “The video begins with a black screen”, contributing to a strong prior in the fine-tuned model.
-We list more qualitative cases and analyses in the Appendix.
+Specifically, InternVL supports up to 13 sub-images, each of which is converted into 256 visual tokens; MiniCPM-8B slices images into a maximum of 10 patches, where each is represented by 96 visual tokens.
+During training and evaluation, we switch off the patchifying option for higher efficiency.
+3.1.2 Training Configurations
+For fairness and ease of reproduction, we follow the official implementations.
+More specifically, we train the whole model end-to-end (except for InternVL-4B, where we freeze the vision encoder) with a learning rate of 5e-6.
 (a) Share-Gemini
@@
 Both datasets mostly cover videos shorter than 3.5 minutes.
-We extract video frames at an FPS of 1.
+We extract video frames at an FPS of 1 for each video..
 For better visibility, we pick samples with frame numbers lower than 99.9 percentile for visualization.
-5.2 The Pitfalls in Fine-tuning
-In light of the limitations of zero-shot inference with image-LLMs, we turn to investigate the fine-tuning approach for improving the video understanding capabilities of models.
-In this section, we first illustrate the training settings and then the experimental findings.
 3.1.3 Training Datasets
-With 8 NVIDIA A800 GPUs, we train end-to-end (except for InternVL-4B, where we freeze the vision encoder) with a learning rate of 5e-6 and a global batch size of 64.
 During our investigation, we utilize two representative types of datasets, i.e., video-caption pairs and video instruction data.
 Specifically, we choose the ShareGemini [32] dataset and the Video-ChatGPT [15] dataset as caption and instruction data, respectively.
-We extract video frames at an FPS of 1 for each video.
-In consideration of efficiency, we use up to 64 frames for InternVL-4B and 24 frames for both MiniCPM-8B and Idefics3-8B.
-When the total number of frames exceeds the thresholds, we uniformly downsample the video frames.
-We show the statistics of video lengths in Fig.
-3 and provide a more detailed introduction of the two datasets below.
+For each video, frames are extracted at an FPS of 1.
+In consideration of efficiency, we use up to 64 frames for InternVL-4B and 24 frames for MiniCPM-8B.
+When the total number of frames exceeds the threshold, we uniformly downsample the video frames.
+The statistics of video lengths are shown in Fig.
+2, and we provide more introduction to the two datasets below.
 ShareGemini-Webvid-core100k.
@@
 The videos in this collection are derived from ActivityNet [36].
-The dataset’s duration is more diverse and uniform, yet the average video length is no more than 3.5 minutes.
+The dataset’s coverage of video duration is larger, yet the average video length is no more than 3.5 minutes.
 There are broadly three types of instructions: video summarization, questions about video content, and creative/generative tasks.
 The dataset is annotated in a semi-automatic manner.
-Some of the data samples are manually annotated by human annotators by refining and enriching the video captions.
+A small portion of data samples are manually annotated by human annotators by refining and enriching the video captions.
 Other instruction data are generated by GPT-3.5 with the aid of off-the-shelf dense prediction and captioning models.
-5.2.2 Low Learning Efficiency of Fine-tuning
-Setting S M L Overall
-Zero-shot 61.3 51.8 44.3 52.5
-30K sampled data 66.2(+4.9) 53.3(+1.5) 47.4(+3.1) 55.7(+3.2)
-200K full data 66.7(+0.5) 54.2(+0.9) 48.1(+0.7) 56.3(+0.6)
-Table 2: Accuracy comparisons of different settings on the Video-MME benchmark, ranking in ascending order regarding overall performance.
-S, M, and L indicate short, medium, and long videos, respectively.
-Zero-shot indicates direct inference with the image-LLM, and full data setting indicates fine-tuning with all the video data samples in the training corpus.
-The performance gains of the latter two settings are calculated relative to the previous row.
-We experiment with fine-tuning with only parts of the video data and with the full volume.
-Specifically, we randomly sample 15% of the total video samples from each video dataset and compare the performance as in Tab.
-2.
-As shown in the table, compared with the short video set, the gains in medium and long video sets are relatively smaller.
-This can be attributed to the fact that the video lengths in the training corpus are mostly shorter than 3.5 minutes, rendering a large gap in the temporal range between training and testing (in the evaluation phase, video lengths range from 4–15 minutes and 30–60 minutes for medium and long videos, respectively).
-More importantly, the additional gain of scaling up training volume from 30K to 200K data samples is limited (+0.6 overall).
-In view of the huge difference in data samples used, the learning efficiency of fine-tuning with these video datasets can be quite limited.
-The phenomenon also suggests that there could be high redundancy in the training corpus.
-5.2.3 A Look at Instruction Diversity
+3.2 Evaluation Setup
+To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [22], MVBench [9], and TempCompass [23] as our benchmarks.
+We do not use traditional video-QA benchmarks (e.g.
+MSVD-QA [19], TGIF-QA [20], ActivityNet-QA [21]) since these benchmarks are generally limited to a small coverage of domains, task types, and video lengths.
+Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel.
+We illustrate more about the benchmarks used as follows:
+Video-MME is a comprehensive benchmark designed for the evaluation of video-LLMs.
+For temporal coverage, videos of short length (up to 2 minutes), medium length (4–15 minutes), and longer duration (30–60 minutes) are included.
+The videos and annotations are manually collected and filtered.
+We only use the raw frames without the subtitles to focus on the evaluation of video understanding capabilities.
+MVBench designs a set of 20 video tasks that cover both perception and cognition, such as scene transition and episodic reasoning.
+Compared to Video-MME, the videos are sourced from existing benchmarks, and the QAs are automatically generated for the 20 pre-defined tasks.
+TempCompass focuses on fine-grained temporal aspects, such as action, speed, and attribute change.
+The videos and meta-information are manually collected, after which annotations are generated by LLMs with the aid of meta-information.
+We use the multiple-choice QA (MCQ) format to align with other benchmarks.
+To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment.
+More details about the implementation of this evaluation scheme are available in Appendix A.1.
+3.3 Main Findings
+3.3.1 Low Learning Efficiency Issue
+Figure 3: Scaling performance of fine-tuning InternVL with different data volumes and types on different benchmarks.
+We consider fine-tuning with ShareGemini, Video-ChatGPT, and a mix of both with a 1:$1:11$ : 1 sampling ratio.
+Training samples are measured in K, where 0 sample indicates zero-shot inference.
 Figure 4: t-SNE plot of instruction distribution of video datasets - ShareGemini and Video-ChatGPT.
 5,000 instructions are sampled from each dataset.
-ShareGemini exhibits some clear clusters, while Video-ChatGPT is relatively more diverse.
-Prior studies have underscored the importance of instruction diversity for fine-tuning LLMs [46] and image-LLMs [42].
-These findings prompt us to conduct an inspection of data in this aspect.
-Thus, we follow previous approaches [36, 45] to visualize the diversity of instructions in the training corpus.
-Specifically, we randomly sample 5,000 instructions from ShareGemini and Video-ChatGPT, respectively.
+The instruction diversity of these two datasets is relatively limited, leading to low learning efficiency.
+Our experiments start with scaling up the training data volume and evaluating the video understanding performance on different general video understanding benchmarks.
+The results are shown in Fig.
+3.
+In general, training either with video caption data (ShareGemini), instruction data (Video-ChatGPT), or a mix of both can boost the image-LLM’s video understanding performance.
+Meanwhile, increasing the training volume brings additional gains in accordance with the data scaling law.
+However, the gains from scaling up quickly reach a plateau.
+For instance, on the Video-MME benchmark, when training with mixed data, 30K samples improve overall accuracy by 3.1 points, while 100K samples only add another 0.5 points, resembling a logarithmic growth.
+In view of this quick and early saturation, the learning efficiency with these video datasets can be quite limited.
+The phenomenon also suggests that there could be high redundancy in the training corpus, and it is possible that we may use less data to achieve a performance comparable to or even better than training with more data samples.
+3.3.2 Probing of Instruction Diversity
+Previous results prompt us to explore the reason for such low learning efficiency.
+Inspired by prior studies, which have underscored the importance of instruction diversity for fine-tuning LLMs [37] and image-LLMs [38], we conduct an inspection of training data in this aspect.
+Specifically, we follow previous approaches [39, 40] to visualize the distribution of instructions in the training corpus.
+5,000 instructions are sampled from ShareGemini and Video-ChatGPT, respectively.
 Then, the instructions are embedded and visualized using the t-SNE technique, as shown in Fig.
 4.
-In the figure, the distribution of ShareGemini exhibits 9 clear clusters.
-This is because the dataset samples from a fixed pool of 9 templates as instructions, each of which is a variant of “Describe this video in detail.” In contrast, the distribution of Video-ChatGPT is more diverse, as it includes not only video summarization but also questions related to specific video content.
-Overall, the instructions for these video samples are not diverse enough.
-6 Our Methods
-Since currently available video data are limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity using synthetic data.
+Overall, the instruction distribution of these two datasets is not diverse enough, which leads to a low data efficiency: The distribution of ShareGemini exhibits 9 clear clusters in the figure, indicating very similar instructions.
+This is because this dataset samples from a fixed pool of 9 templates as instructions, each of which is a variant of “Describe this video in detail”.
+On the other hand, the distribution of Video-ChatGPT seems relatively more diverse, as it includes specific questions related to video content and details besides video summarization.
+Nevertheless, the instruction diversity is still low due to the nature of self-instruction and a few fixed task-specific prompting templates for data curation.
+4 Methods
+4.1 Design Concept
 Figure 5: A conceptual illustration of our scheme for synthesizing video-like samples from text data.
@@
 Zoom in for better visibility.
-A rich source of instruction data lies in the text domain, and they can effectively enrich the diversity of instruction.
-Nevertheless, there is inherently a modality gap between the text and visual domain.
+Since currently available video data can be limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity by incorporating new synthetic data.
+A rich source of instruction data lies in the text domain, and they can effectively complement the vision domain.
+Nevertheless, there is inherently a modality gap between the text and visual domains.
 To better utilize these data, we bridge the modality gap by synthesizing images with the text.
 Fig.
-5 illustrates our overall idea.
-Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments and embed the text of each segment into a sequence of images.
-For example, the context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g.
-an inquiry to summarize the given paragraphs and the corresponding synopsis.
-After the transformation, the structure of these instruction data is exactly the same as the video counterparts, and we can incorporate the synthetic data into the video training corpus without extra processing.
+5 illustrates our overall workflow for synthesis of data.
 Figure 6: t-SNE plot of instruction distribution after applying our proposed methods.
@@
 LongAlpaca [41] and LongQLoRA [41], the overall distribution becomes more diverse.
-5,000 instructions are sampled from each dataset.
 Our proposed scheme enjoys three benefits: (1) Mixing in text data can effectively enrich the instruction diversity (Fig.
@@
 Thus, utilizing synthetic data can be economical.
-On a related note, LLaMA-VID [20] and LongVA [43] also utilize text data to assist video understanding, and we underscore the differences here: (1) Different objectives: LLaMA-VID and LongVA use text data in the fine-tuning and continued pre-training stage, respectively, in order to expand the context window for long video understanding.
-In contrast, our methods mainly utilize the diversity of instructions in the text data to facilitate efficient fine-tuning.
-(2) Different processing methods: LLaMA-VID and LongVA directly use the pure text data in training, while our methods transform text samples into video-like samples and unify the training format.
+4.2 Implementation Details
+Each text sample is a (long-context, question, answer) triplet.
+For example, the long context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g.
+an inquiry to give a synopsis, or questions related to the paper.
+After the data transformation process, each sample is a video-like (images, question, answer) triplet, where long-context information is transformed into a series of images, and the question and answer stay unchanged.
+The key to the data synthesis procedure is synthesizing images with pure text.
+Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments according to word counts (set to 115 empirically) using an open-sourced NLP toolkit [^2] https://www.nltk.org/.
+These text chunks are then transformed into a sequence of images separately.
+Specifically, each text chunk is embedded into a blank image with a white background.
+This is achieved using bitmap font with the ImageFont module of Pillow library [^3] https://pillow.readthedocs.io/en/stable/.
+Each image is 448x448 pixels in size, and the font is 20 pt large, black color, Arial Regular type.
+We use a bounding box to control the layout, leaving a margin of 20 pixels on each side, so each line of the text has roughly the same width.
+After the transformation, the structure of these synthesized data is exactly the same as that of video samples, and we can seamlessly incorporate the synthetic data into the video training corpus.
+5 Evaluation on Proposed Methods
+This section includes experimental results and discussions of our proposed method, including (1) a comparison with mainstream methods, (2) an ablation study on data mixes, and (3) an examination of key properties, including data scaling performance and gains in long video understanding.
+Methods Size Frames Short Medium Long Overall
+Proprietary Models
+GPT-4V [42] N/A 10 70.5 55.8 53.5 59.9
+Claude 3.5 Sonnet [43] N/A 20 71.0 57.4 51.2 60.0
+GPT-4o [44] N/A 384 80.0 70.3 65.3 71.9
+Gemini 1.5 Pro [45] N/A 1fps 81.7 74.3 67.4 75.0
+Open-Source Models
+VideoChat2 [9] 7B 16 48.3 37.0 33.2 39.5
+Video-LLaVA [13] 7B 8 45.3 38.0 36.2 39.9
+Chat-UniVi-v1.5 [14] 7B 64 45.7 40.3 35.8 40.6
+VideoLLaMA 2 [8] 7B 16 56.0 45.4 42.1 47.9
+VITA [46] 8x7B 32 65.9 52.9 48.6 55.8
+Kangaroo [47] 8B 64 66.1 55.3 46.6 56.0
+VITA-1.5 [48] 7B 16 67.0 54.2 47.1 56.1
+FT w/ InternVL [7]
+Zero-shot 3.8B 64 61.3 51.8 44.3 52.5
+200K video data 3.8B 64 66.7 54.2 48.1 56.3
+Sparrow (30K hybrid data) 3.8B 64 67.0 53.7 49.3 56.7
+Table 1: Accuracy comparisons of different methods on the Video-MME benchmark.
+Performance is ranked in ascending order regarding overall performance (The models are grouped according to open-source or not).
+Our method uses only 15% of the total sample size compared to the full volume video datasets (200K) for fine-tuning and achieves comparable performance.
+30K hybrid data comprise 20K data sampled from video datasets and 10K synthesized from our method.
+Bold digits indicate the best performance within each group.
+5.1 Comparison with Mainstream Methods
+We compare the results with some representative proprietary models, including GPT-4V [42], Claude 3.5 Sonnet [43], GPT-4o [44], Gemini 1.5 Pro [45], and open-source video-LLMs of similar LLM parameter size, including Video-LLaVA [13], VideoChat2 [9], Chat-UniVi-v1.5 [14], VideoLLaMA 2 [8], VITA [46], and Kangaroo [47], as summarized in Tab.
+1.
+The table results show that, through zero-shot inference, the image-LLM Intern-VL already outperforms a variety of video-LLMs with larger LLM parameter sizes.
+This might be due to the rich pre-trained knowledge embedded in the model parameters since the image-LLM has been trained with large-scale and high-quality image-text data.
+The vision prior lays a strong foundation for further video fine-tuning, where models learn temporal and causal concepts from activities, events, etc.
+The model fine-tuned with full video datasets achieves an overall gain of 3.8 points on the image-LLM, closing the gap between open-source models and proprietary ones.
+Notably, our methods use only 15% of the total sample size compared to the full volume (200K) for fine-tuning and achieve comparable performance.
+This result suggests the high data efficiency of our proposed scheme since mixing in synthetic samples mitigates the low instruction diversity issue illustrated in the earlier section.
+5.2 Ablation on Different Data Compositions
 Data Mix S M L Overall
@@
 200K full data 66.7 54.2 48.1 56.3
-Table 3: Accuracy comparisons of different data compositions on the Video-MME benchmark.
+Table 2: Results of different data compositions on the Video-MME benchmark.
 Our proposed scheme achieves an overall performance superior to other data mixes of the same amount (30K) and even more data (200K).
 The Zero-shot and 200K full data fine-tuning settings are listed for comparison.
-Video-MME MVBench TempCompass Avg.
-Short Medium Long Overall
-Text-aided Models
-LLaMA-VID [20] 40.6 40.2 34.9 38.6 41.9 45.6 42.0
-LongVA [43] 61.1 48.8 45.4 51.8 50.7 56.9 53.1
-MiniCPM-8B [37]
-Zero-shot 58.3 46.1 40.1 48.2 42.9 49.1 46.7
-200K data 60.3 47.3 44.7 50.8 48.0 54.7 51.2
-30K data 61.9 51.7 45.6 53.0 48.4 56.8 52.7
-Idefics3-8B [17]
-Zero-shot 61.3 47.0 45.1 51.2 49.6 55.9 52.2
-200K data 62.0 53.1 44.9 53.3 50.7 62.9 55.6
-30K data 67.1 53.8 48.0 56.3 51.6 62.3 56.7
-Table 4: The performance of different models/settings on three video benchmarks.
-Our methods use only 15% of the total sample size compared to the full video sets (200K) for fine-tuning and achieve similar or even superior performance.
-The highest performance of a model among all settings is highlighted in bold.
-LLaMA-VID and LongVA also use pure text data but without transformation, which is different from our methods and objectives.
-7 Evaluation on Proposed Methods
-7.1 Ablation on Different Data Compositions
 In order to examine the impact of different data compositions and validate the effectiveness of the proposed method, we conduct an ablation study and construct the following settings with the same amount of total data samples:
@@
   • Same video samples as above (20K in total), plus 10K samples of corresponding pure text data.
-The effectiveness of our proposed methods.
-As shown in Tab.
-3, comparing the first three rows, we can find that when using the same amount of video samples, training only with ShareGemini is not as effective as using more diverse data compositions.
-Meanwhile, using the same amount of data, our proposed scheme (the 4th row) achieves the best performance.
-Moreover, when compared with the full 200K data fine-tuning setting, our proposed scheme uses much fewer data samples (only 15%) to achieve comparable performance, and the training cost reduces from 276.8 GPU hours to 33.6 GPU hours, making an 8.2x speedup.
-The overall results demonstrate the importance of instruction diversity and the effectiveness of our proposed methods.
+[…]
 Notably, replacing the synthetic data with the original pure text counterpart achieves an overall inferior performance.
 We hypothesize that this is due to the inherent domain gap between vision and text.
-Thus, to simulate the structure of video sequences, transcribing long text into images might be necessary.
+Thus, to simulate the structure of video frame sequences, transcribing long text into images is necessary.
+Figure 7: Scaling performance with different data volumes and types on general video benchmarks.
+We consider four settings, i.e.
+fine-tuning with video data only, our hybrid data (video data mixed with our synthetic samples with a ratio of 2:1:212:12 : 1), TOPA data, and T3 data.
+By default, video data (the caption dataset and the instruction dataset) are mixed with a 1:$1:11$ : 1 sampling ratio.
+Training samples are measured in K, where 0 sample indicates zero-shot inference.
+Purple-shaded area indicates gains of our methods over the video data baseline.
 Can synthetic data help models understand longer videos?
@@
 However, on the long video benchmark set, our proposed method still achieves a score that is 1.2 points higher than the full data training (as shown in the 4th row compared to 200K full data in Tab.
-3).
+2).
 This result suggests that fine-tuning with a long multimodal context can enhance the comprehension of longer videos.
-In the following section, we will present additional results to further illustrate this point.
-7.2 Generalization Ability of Our Proposed Scheme
-We further verify the generalization ability of the proposed scheme and evaluate our methods on more image-LLMs of larger parameter sizes, including MiniCPM-8B and Idefics3-8B, across different benchmarks, as shown in Tab.
-4.
-General effectiveness of the proposed scheme.
-The results in the table suggest that on both models, though fine-tuning with fewer samples, our proposed scheme achieves performance on par with or better than fine-tuning with the complete video datasets.
-We also compare with related methods, LLaMA-VID and LongVA, which also utilize text data but in the original form of pure text, in order to expand the context window.
-In contrast, our methods excel in enriching the instructions via synthesizing video-like data from text data.
-Boost in long video understanding.
-On the long video sets of the Video-MME benchmark, our methods provide consistent gains compared to the full data fine-tuning setting, though we use video samples (most no longer than 3.5 minutes in the training sets) of lengths much shorter than the test sets (30-60 minutes).
-The results indicate that our proposed methods can assist models in understanding long videos.
+In the following section, we will present additional results and discussions to illustrate this point further.
+[…]
+Baseline 0 24 40.1 40.0 44.5 41.6
+30 24 44.7 39.7 45.4 43.3
+60 24 46.2 42.7 46.2 45.1
+100 24 46.7 44.1 45.3 45.3
+Our Method 30 24 45.6(+0.9) 48.7(+9.0) 51.4(+6.0) 48.5(+5.2)
+60 24 46.2 51.2(+8.5) 53.2(+7.0) 50.2(+5.1)
+100 24 48.7(+2.0) 50.1(+6.0) 57.0(+11.7) 51.9(+6.6)
+Table 3: Long video understanding performance with different data volumes and types.
+The Baseline method adopts only video data, while Our Method utilizes a mix of video data and synthetic samples attained from our method.
+In the Baseline group, 0 sample indicates zero-shot inference.
+The performance gains are calculated relative to the video data fine-tuning baseline trained with the same number of samples.
+5.3 Examination of Key Properties
+In this section, we further examine key properties of our proposed method, including general effectiveness, scaling performance, and effectiveness when adopted in long video understanding scenarios.
+5.3.1 General Effectiveness and Scaling Performance
+We further verify the proposed scheme’s effectiveness by evaluating our methods on another image-LLM of larger parameter size, i.e.
+MiniCPM-8B, across different benchmarks.
+Specifically, through scaling up with different volumes and types of data, we compare our methods against a pure video data baseline, as well as other relevant methods, including TOPA and T3.
+Both TOPA and T3 first translate vision information into text, such as captions and relations between objects.
+Then, synthetic text QAs are constructed to simulate video reasoning samples, aiming to transfer temporal reasoning capabilities from text to video.
+Note that since the original data format of TOPA is largely different from the current paradigm, we design a template to adapt them to the instruction data format (More details are available in Appendix A.2).
+The results are summarized in Fig.
+7.
+General effectiveness.
+When using the same amount of training samples, our methods almost always outperform other methods by a clear margin in all the evaluated benchmarks.
+Specifically, when using 30K samples, our method achieves an overall accuracy of 52.7, surpassing the baseline by 3.9 points.
+Notably, it even outperforms the baseline trained with 100K samples by 1.7 points.
+Similarly, on the MVBench benchmark, with 100K samples, our method attains a 4.3-point absolute gain over the baseline.
+Overall, the superior performance on different benchmarks showcases the general effectiveness of our proposed method.
+Scaling performance.
+A clear issue with other methods is that they are more prone to performance saturation when scaling up the data budget.
+For instance, the baseline method uses 60K samples to improve the Video-MME benchmark by 3.1 points, while using 100K samples only achieves another 0.9 points of absolute gains.
+In contrast, our proposed scheme shows more stable and consistent improvements when scaling up the data volumes compared with other methods.
+This suggests that ensuring a more diverse distribution of instruction in video training is very important, or otherwise, the learning efficiency may downgrade.
+To avoid such circumstances, the proposed data augmentation method can be an effective approach.
+Discussion: Can we scale up only with synthetic textual samples?
+An intriguing and highly relevant question is whether we can scale up the synthetic samples without using any real video samples.
+Since the text is more compact and less redundant than a whole video, training in this way is more economical.
+Unfortunately, the empirical results show that this is probably unfeasible.
+As shown in Fig.
+7, scaling with synthetic text data (TOPA and T3) shows undesirable characteristics, i.e.
+this approach can easily reach the saturation point or even slightly downgrade.
+Other critical issues include the modality gap and special processing of videos in various domains (such as egocentric videos and movies).
+Besides, since text suffers inevitably from information loss when translated from videos, text data might be better used as a supplement to videos, which helps inject temporal reasoning language prior into the LLM backbones (similar to TOPA or T3) or mixing with video data as a regularization method (as our method does).
+Figure 8: Long video understanding performance evaluated with a different number of frames.Video context” denotes the max number of video frames used in training, i.e.
+24 frames; “LLM context” denotes the context window of the LLM backbone.
+Purple-shaded area indicates performance gains of our proposed method over the video data baseline.
+5.3.2 Long Video Understanding Performance
+We adopt tailored benchmarks to evaluate long video understanding capabilities, including LongVideoBench [49], MLVU-M [50], and the long video set of Video-MME, and report the performance on evaluation sets for the former two benchmarks.
+Our study focuses on two aspects: (1) performance improvement in terms of long video understanding compared with the video fine-tuning baseline and (2) frame number (multimodal context) generalization ability in the inference stage.
+Performance improvement w.r.t.
+training sample size.
+As shown in Tab.
+3, we observe consistent improvements in long video understanding compared with the video fine-tuning baseline despite an absence of any long video training data.
+Notably, when using the same number of 100K training samples, our hybrid data scheme improves by 6.6 points over the baseline.
+We hypothesize this is because the reasoning patterns within the long text can also transfer to video understanding, which exempts the need to curate specific text samples translated from vision information [27, 28].
+Performance change w.r.t.
+frame number.
+We also investigate if our approach expands the context window.
+Larger context windows are usually beneficial since video-LLMs derived from large-context LLM backbones often accompany frame number generalization and benefit from inputting more video frames [49, 26].
+However, we do not observe this trend.
+As shown in Fig.
+8, the model fine-tuned with long synthetic samples still follows a similar pattern, that is, when performing inference beyond the video context of the training stage, the performance stays relatively stable and does not benefit from more frame input.
+And when the frame number exceeds the LLM context, the performance plunges to a low level.
+Thus, promising directions for further improvements may include continued pre-training to expand the LLM context window.
 6 Conclusion
-This paper revisits the two mainstream approaches of utilizing image-LLMs to perform video understanding.
-A chain of investigations reveals possible drawbacks in both approaches, underscoring the bottlenecks of zero-shot inference and the easily ignored aspects of fine-tuning.
-Among these discoveries, the limited instruction diversity of current video data is notable, leading to downgraded learning efficiency.
-Aimed at this specific issue, we develop an efficient method, dubbed T2Vid, to enrich the instructions in the training corpus.
-Comprehensive evaluations demonstrate the effectiveness of the proposed method.
-We hope the findings from this paper can spark deeper thinking about using MLLMs for video understanding and how to curate video data of higher quality.
-Appendix A Implementation Details
+In this paper, we propose Sparrow, a data-efficient training scheme for video-LLMs, which enables training with fewer samples and achieving better video understanding performance.
+This method derives from our empirical findings that the low learning efficiency in data scaling may be ascribed to a limited instruction diversity in the training corpus.
+Thus, we design an economical data augmentation method that synthesizes video-like samples rich in instruction diversity.
+Comprehensive experiments demonstrate the general effectiveness and key properties of our proposed method.
+We hope this paper’s findings can spark more explorations of efficient training and high-quality video training corpora.
+Appendix A More Implementation Details
 A.1 Answer Judging
@@
 Our judging prompt for the LLM is modified from VLMEvalKit [52], as shown in Tab.
-5.
-Appendix B Additional Results in Experiments
-B.1 Results on Different Frame Counts
-Figure 7: Accuracy of MiniCPM-8B and Idefics3-8B on the Video-MME benchmark when inference with different numbers of video frames.
-As input frames increase, the performance of both models stays relatively stable and does not improve much.
-“max training frames” denotes the max number of patches used for a single image when training; “max training context” denotes the maximum context window set when training; “max LLM context” denotes the context window of the original LLM backbone before further multimodal training.
-We present the performance of different input frame numbers of MiniCPM-8B [37] and Idefics3-8B [17] in Fig.
-7.
-As input frames increase, the performance within the context window of the LLM backbone is more stable than the InterVL-4B, which is of a smaller parameter size.
-Nevertheless, the performance still does not improve much as more frames are available, again indicating that the effective context is limited.
-Thus, simply expanding the LLM context window for image-LLMs to facilitate long video understanding would not be very practical.
-As in the main paper, we attribute this phenomenon to the lack of capabilities of image-LLMs to truly understand the temporal relationships among frames.
-Additional case studies related to this topic are available in the following section.
-B.2 Case Study
-In the main paper, we identify the drawback of image-LLMs generally failing to understand the temporal relationships among video frames.
-In this section, we supplement more cases and analyses on this topic.
-Figure 8: A case study using the prompt “Describe this video.”, where the video is a documentary of the Arctic environment.
-Compared with the fine-tuned model, the zero-shot case reasons with and describes video frames without understanding the temporal correlations.
-This pattern follows the common behavior of the image model.
-The keyframes are highlighted with colors (except for red, which is used to indicate hallucinations).
-The numbers on the upper-left corner of each frame indicate frame order.
-Zoom in for better visibility.
-In case 1 (Fig.
-8), we list a case where we instruct models to describe the video, while the image-LLM fails to understand the separate scenes and temporal relationships but instead focuses on several specific frames and describes rather independently.
-In contrast, the model fine-tuned with video data can more accurately distinguish the actions happening in the video and understand the connections in a temporally coherent way.
-Figure 9: A case study designed to test the temporal reasoning capabilities of models.
-In the video, three polar bears show up at the beginning, and in the end, the scene changes into a video shot of two polar bears.
-The image-LLM fails to recognize the change through zero-shot prompting.
-Note that we have verified that the image-LLM can correctly recognize the number by inputting a single frame separately.
-In case 2 (Fig.
-9), we ask the two models to identify the temporal changes in object counts.
-The fine-tuned model correctly recognizes the changes in the number of polar bears as scenes change in the video clip, while the image-LLM fails to do so and seems to give the answer according to the last few frames.
-This is expected since image-LLMs are trained with static images rather than dynamic videos, where the latter inject temporal concepts into models.
+4.
+A.2 Reproduction Details of Baseline Methods
+Due to an inconsistent formulation between the method TOPA [27] and our proposed method, we adapt the implementation for a fair comparison.
+The original sample comprises a global caption for the whole video and frame-specific information.
+The frame-related information contains a frame-level caption and some descriptions of key objects in the frame.
+Thus, we design a prompt template to fit the original textual samples into the unified training format.
+A real case of formatting the sample with the devised template is shown in Tab. 5.
 System message
@@
 You are provided with a question, several options, and an answer, and you need to find which option is most similar to the answer.
-If the meaning of all options are significantly different from the answer, output Z.
-Your should directly output a single uppercase character, such as A, B, C, D (if they are valid options) and Z, and nothing else.
+If the meaning of all options is significantly different from the answer, output Z.
+You should directly output a single uppercase character, such as A, B, C, D (if they are valid options) and Z, and nothing else.
 Here are two examples.
 Example 1:
-Question: What is the main object in image?
+Question: What is the main object in the image?
 Options: A. teddy bear.
@@
 Example 2:
-Question: What is the main object in image?
+Question: What is the main object in the image?
 Options: A. teddy bear.
@@
 Output: Z
-Now here is the question, options, and the answer, you should match and give me the option letter:
+Now here are the question, options, and the answer, you should match and give me the option letter:
 Question: {Question}
@@
 {Model Answer} is the raw prediction of MLLMs.
+User:
+You will be provided with some information about a video, including a global caption for the whole video, global captions for each video frame, and descriptions of key objects in each frame.
+Answer the questions using the information below.
+Video information:
+Caption: This video shows the carpentry process.
+At first, the person sits on a workbench and measures the length of the plank.
+Then, he uses a saw to cut the wooden plank into multiple pieces.
+Then, he uses a hammer to nail two pieces of wood together.
+Finally, he takes a break and drinks water, and leaves the workshop.
+Frame information:
+Frame 1:
+Caption: the person sits on a workbench and holds a hammer in his right hand.
+the person: the person is sitting on a workbench in the center of the frame. He is looking at a wooden plank in front of him.
+hammer: He has a hammer in his right hand.
+workbench: He is sitting on a workbench in the center of the frame.
+Frame 2:
+Caption: the person takes a chisel from a red toolbox and starts chiseling the plank.
+chisel: There are a chisel and a hammer on the workbench in the center of the frame.
+the person: the person is holding a chisel in the center of the frame.
+red toolbox: There is a red toolbox in the right corner of the frame.
+workbench: He is sitting on a workbench in the center of the frame.
+Frame 3:
+Caption: the person measures the length of the plank with a measuring tape.
+the person: the person carries a pencil and a ruler in his right hand.
+pencil: There is a pencil in the right hand of the person.
+measuring tape: the person is holding a measuring tape in his right hand.
+wooden plank: The plank is above the workbench.
+Frame 8:
+Caption: the person puts his hammer and chisel into the toolbox and leaves the workshop.
+toolbox: the person is putting a hammer and a chisel into a toolbox in the center of the frame.
+the person: the person is walking toward the door in the center of the frame.
+hammer: There are a hammer and a chisel in the right hand of the person.
+workshop: the person is leaving the workshop.
+Question: Why does the person use a measuring tape in the third frame?
+A. To hammer the plank
+B. To cut the plank
+C. To measure the length of the plank
+D. To handle the saw
+E. To fix the length of the wooden plank
+Answer with the option’s letter from the given choices directly.
+Assistant:
+C
+Table 5: An example of using the template for structuring textual samples into the training format. For simplicity, we only show part of the frame-level information.